MEASURING THE QUENCHING OF NO FLUORESCENCE PRODUCED FROM THE EXCITATION OF PHOTO-FRAGMENTED NITROBENZENE USING A PICOSECOND LASER.

CHRISTOPHER J. LUE, CHAKREE TANJAROON, \underline{J. BRUCE JOHNSON}, SCOTT W. REEVE, Arkansas Center for Laser Applications and Science and Department of Chemistry and Physics, P.O. Box 419 State University, AR 72467; SUSAN D. ALLEN, Embry Riddle Aeronautical University, 600 S. Clyde Morris Boulevard, Daytona Beach, FL 32114.

The military is interested in using spectroscopic methods to detect nitroaromatic compounds related to explosives. Upon absorption of a UV photon, nitrobenzene can dissociate into C_{6}H_{5}O and NO. Wynn, et al. ^{a} have shown that looking at NO fluorescence from the photodissociated nitrobenzene could be a possible detection method. However, the fluorescence can easily be quenched by molecular oxygen and other constituents in air. We have measured fluorescence lifetimes of the nascent NO resulting from photo-fragmented nitrobenzene using a pulsed picosecond tunable laser (pulse width ≈15 ps) by means of a two-color process. In the two-color process, photons of a particular energy dissociated the nitrobenzene while photons of a different energy probed the A^{2}Σ^{+} ← X^{2}Π_{(1/2,3/2)} NO band system between 225-260 nm. We have performed the measurements with different background pressures of He, N_{2}, and air. We present the results of these measurements which indicate considerable quenching of the NO fluorescence due to oxygen.

aWynn, C. M.; Palmacci, S.; Kunz, R. R.; and Rothschild, M.Opt. Express, OSA, 2010, 18, 5399-5406